Click Here to receive Service Notes automatically via email.
|07/05/2002||On Monday evening, 07.08.02, the main administrative webserver, Argon, experienced a major hardware failure. Web services were unavailable for several hours while temporary repairs were made. IBM has been contacted and will be on-site Wednesday (07.10.02) to make further repairs. Web services may be intermittently unavailable during this time.|
|07/05/2002||We are currently experiencing difficulties between the Truman and AOL email systems. Email sent from Truman users to AOL users is not being delivered. We are working with AOL to resolve this issue as soon as possible. The AOL estimate of when the problem will be resolved is by Monday morning August 5th|
|07/11/2002||1:00pm, Friday 8.2.02 - Truman's connection to the Internet is down. Technicians are currently investigating the problem. All Truman websites should continue to be available.
|08/04/2002||A problem with Truman's firewall caused the temporary disruption in our connection to the Internet off-campus. The firewall was rebooted and service was restored by 2pm.|
|08/07/2002||Difficulties between the Truman and AOL e-mail systems have been resolved for now. E-mail sent to AOL users should now be processed properly.|
|08/11/2002||MOREnet has announced a potential period of service interruption as they perform upgrade maintenance in St. Louis over the weekend. Steps have been taken to minimize network interruptions, however connectivity may be interrupted between 5:30 and 6:00pm tonight (Friday, 8.9.02) and again on Sunday (8.11.02) from 10:00 to 10:30am. Actual downtimes are expected to be brief.|
|07/25/2002||IBM has completed work on the administrative webserver, Argon. All sites are back up and running. Web services were unavailable from approximately 12pm to 1pm.|
|07/23/2002||On Monday evening, 07.08.02, the main administrative webserver, Argon, experienced a major hardware failure. Web services were unavailable for several hours while temporary repairs were made. Services on Argon are back up and available, but the system is running without one CPU.
IBM has been contacted and will be on-site tomorrow (07.10.02) to make further repairs. Web services may be intermittently unavailable during this time.
Post Service Note